OpenStudy (anonymous):

given that π/2 < or equal to θ < or equal to 3π/2 and tan θ=3, find the value of cos 2θ

OpenStudy (dumbcow):

 \[\cos(2\theta) = \cos^{2} \theta - \sin^{2} \theta\] \[\cos(2\theta) = \frac{1}{10}-\frac{9}{10} = -\frac{8}{10}\]